OpenStudy (anonymous):

weight of a body of mass m is decreased by 1% when it is raised to height h above the earth surface. if the body is taken to a depth h in a mine,then its weight will be??

OpenStudy (anonymous):

sry its g not G

OpenStudy (anonymous):

ans is 0.5% change

OpenStudy (anonymous):

it is said it is 0.5%decreases i think u didnt read ques again by concn. as here it is said abu weight plz read it and explain me

OpenStudy (jamesj):

The weight of a body of mass m in the gravitational field of the earth is \[ W = \frac{GMm}{r^2} \] where M is the mass of the earth and r is the radial distance of the object from the center of the earth If the mass is at a height h over the surface of the earth, then its distance from the center of the earth is r = R + h where R is the radius of the earth. The weight of an object on the earth is \[ W = \frac{GMm}{R^2} \] If at height h, i.e., r = R + h, it is 1% less, that means \[ W = \frac{GMm}{r^2} = 0.99 \frac{GMm}{R^2} \] Re-arranging this equation, this means that \[ r^2 = (R+h)^2 = (1/0.99) R^2 \] i.e., \[ R + h = \sqrt{1/0.99} R \] i.e., \[ h = (\sqrt{1/0.99} - 1)R -- (*)\] Now, if you're in a mine at a depth of h, the radial distance becomes \[ r = R - h \] Using the expression (*) for h, \[ r = R - h = (2 - \sqrt{1/0.99})R \] Calculating this, to five decimal places \[ r = 0.99496R \] Hence \[ r^2 = 0.9899 R^2 \] and \[ 1/r^2 = 1.01016 1/R^2 \] and at this depth of r = R - h, the weight of the mass m is \[ \frac{GMm}{r^2} = 1.01016 \frac{GMm}{r^2} \] In other words, the mass is 1.016% more.

OpenStudy (jamesj):

Correction of last equation: \[\frac{GMm}{r^2} = 1.01016 \frac{GMm}{R^2} \]

OpenStudy (anonymous):

:( the ans is 0.5% decreases from its intital wieght

OpenStudy (anonymous):

thanX for afforts but i ll ask my teacher now i dont think i can get ans.. but thnx for ur help :)